Calculate the values of the six trigonometric functions of the angle θ given in standard position, if the terminal side of θ lies on the given line.

y=8x, x≥0

Answer by Alan3354(69443)

Calculate the values of the six trigonometric functions of the angle θ given in standard position, if the terminal side of θ lies on the given line.
---------
Use the point (1,8)
---
x = 1
y = 8
r = sqrt(65)
----
sin = y/r
cos = x/r
tan = y/x
cot = x/y
sec = r/x
csc = r/y